# Exploring Efficient Open-Vocabulary Segmentation in the Remote Sensing

### v1

@@ Appendix A A: More related works
Semantic segmentation models traditionally rely on predefined datasets, limiting their scalability and adaptability to novel scenarios. The advent of pretrained Vision-Language Models (VLMs) has introduced a new paradigm for segmentation. However, these models, pretrained at the image level, lack the fine-grained pixel-text alignment required for effective segmentation, posing a significant challenge for direct application. To address this issue, researchers have explored two main strategies: two-stage and single-stage approaches. Two-stage methods (Liang et al. 2023; Xu et al. 2022; Zhou, Loy, and Dai 2022) utilize external mask generators to refine segmentation predictions. However, the limited accuracy of these generated masks reduces the overall effectiveness and generalization of such approaches. In contrast, single-stage models integrate mask generation directly within the segmentation process. FC-CLIP (Yu et al. 2023) and its enhanced version (Jiao et al. 2025) incorporate trainable mask predictors within the CLIP framework, improving segmentation accuracy. More advanced techniques, such as Cat-Seg (Cho et al. 2024) and SED (Xie et al. 2024b), further refine segmentation performance by leveraging similarity matrices as pseudo-masks. Recent advancements have introduced novel mechanisms to enhance segmentation effectiveness. Side adapters, proposed in (Xu et al. 2023), facilitate efficient feature fusion, while frequency-domain modules (Xu et al. 2024) improve model generalization. Additionally, (Shan et al. 2024) presents an adaptive integration of SAM and CLIP outputs, optimizing segmentation efficiency. These ongoing innovations continue to refine vision-language alignment, advancing both accuracy and adaptability in open-vocabulary segmentation.

### v2

@@ Appendix A A: More related works
Semantic segmentation models traditionally rely on predefined datasets, limiting their scalability and adaptability to novel scenarios. The advent of pretrained Vision-Language Models (VLMs) has introduced a new paradigm for segmentation. However, these models, pretrained at the image level, lack the fine-grained pixel-text alignment required for effective segmentation, posing a significant challenge for direct application. To address this issue, researchers have explored two main strategies: two-stage and single-stage approaches. Two-stage methods (liang2023open; xu2022simple; zhou2022extract) utilize external mask generators to refine segmentation predictions. However, the limited accuracy of these generated masks reduces the overall effectiveness and generalization of such approaches. In contrast, single-stage models integrate mask generation directly within the segmentation process. FC-CLIP (yu2023convolutions) and its enhanced version (jiao2025collaborative) incorporate trainable mask predictors within the CLIP framework, improving segmentation accuracy. More advanced techniques, such as Cat-Seg (cho2024cat) and SED (xie2024sed), further refine segmentation performance by leveraging similarity matrices as pseudo-masks. Recent advancements have introduced novel mechanisms to enhance segmentation effectiveness. Side adapters, proposed in (xu2023side), facilitate efficient feature fusion, while frequency-domain modules (xu2024generalization) improve model generalization. Additionally, (shan2024open) presents an adaptive integration of SAM and CLIP outputs, optimizing segmentation efficiency. These ongoing innovations continue to refine vision-language alignment, advancing both accuracy and adaptability in open-vocabulary segmentation.